Two vectors , vector A and vector B ,make angles40°and 120° respectively with the X axis.If magnitudeof vector A is 6 unit and that of B is 5 unit.Find the resultant of these vectors.

Grade:11

1 Answers

Anmol
6 years ago
Angle between vectors A and B=120-40=80°.Resultant= sqrt(a^2+b^2+2abcos(theta))=sqrt(61+60cos(80°))